Ariel had a fruit basket with 24 oranges and bananas. She had 3 times more oranges than bananas. She dropped 2 times more oranges than bananas. There were 15 pieces of fruit left.

Answer by richwmiller(17219)

24(xy)+3x-2y=15
huh??
o+b=24
3b=o
24-x=15
x=9
O+B=x
2B=O